Prove: A x B = ∅ iff.  A = ∅ or B = ∅
Answer #1

To show & Ax B= 0 + A=0 or B= 6 (6 let Ax B=0 On the contrary let us assume that A=0 and B+0 Ax B = (1,4) 11e A, YEBY A there